# Learning Chaotic Dynamics with Neuromorphic Network Dynamics

### v1

@@ IV Conclusions
These results validate the hypothesis that the internal and network dynamics of memristive networks can be used in a reservoir computing approach to learn complex chaotic dynamics. It was found that the network dynamics can be optimised for predicting the Lorenz attractor by appropriate scaling of the external input voltage to values that are not too low or too high. Optimal intermediate values of input voltage scaling were found to drive internal memristive states that traversed the entire dynamical range, from low to high conductance states, thereby maximising use of available internal degrees of freedom. Interestingly, this study also revealed that some nonlinear dynamical features, such as high frequency fluctuations, were effectively filtered out by ridge regression in the external output layer. This suggests that physical reservoir computing may be robust against noise and variations due to the specific internal dynamics of the memristive material system. On the other hand, our results also indicate that it is possible to suppress such features by ensuring electrodes provide sufficient coverage of network nodes to sustain dynamic activity across the memristive edges.
@@ Appendix B Nonlinear Responses in Memristive Edges
Figure A.1: (a) Total dynamics measure (i.e. $\delta_{\rm low,med}^{i,j}+\delta_{\rm med,high}^{i,j}+\delta_{\rm low,high}^{i,j}$) and (b) nonlinearity measure $m_{\rm nonlin}^{i,j}$ with respect to the maximum and minimum of each corresponding memristive edge voltage signal across time. The shaded regions are where edge voltage signal amplitudes can cross $V_{\rm set}$ and $V_{\rm reset}$, i.e. where $\dot{x}_{ij}$ may be nonzero. All data points are obtained from the same experiment for $\alpha=0.2\,{\rm V}$ in Fig. 11. Edges connected to the input nodes are omitted, and the figures are zoomed in for clarity.
@@
Figure A.2: By artificially inducing a highly nonlinear (top, red) $x$ signal into the edge between node 2 and 1 of this simple 3-node toy model, unwanted nonlinear features (bottom, red) can be observed in the readouts of node 1, which are vastly different from the original (orange) signal. This also affect the $x$ (top, blue) signals of the edge between node 1 and 0. These memristive edges uses the tunnelling model.
@@
Figure A.3: By artificially inducing a highly nonlinear (top, red) $x$ signal into the edge between node 2 and 1 of this simple 3-node toy model, unwanted nonlinear features (bottom, red) can be observed in the readouts of node 1, which are vastly different from the original (orange) signal. This also affect the $x$ (top, blue) signals of the edge between node 1 and 0. Unlike Fig. A.2, these memristive edges uses the binary model.
@@
From Eq. (10) it is evident that $x$ will only increase in amplitude when the voltage amplitude at the corresponding memristive edge crosses $V_{\rm set}$, and decrease in amplitude when it crosses $V_{\rm reset}$. This will occur in two scenarios: if the maximum edge voltage signal (across time) of the edge between node $i$ and node $j$ satisfies $\max_{t}(v_{ij}(t))>V_{\rm set}$, while the minimum edge voltage satisfies $\min_{t}(v_{ij}(t))<V_{\rm reset}$, or $\max_{t}(v_{ij}(t))>-V_{\rm reset}$ and $\min_{t}(v_{ij}(t))<-V_{\rm set}$. This results in two overlapping regions where dynamical activity can be observed. Indeed as seen in Fig. A.1(a), the dynamic measure is only nonzero within this specified (shaded) region, and it is clear that larger ranges (i.e. larger $\max_{t}(v_{ij}(t))$ and smaller $\min_{t}(v_{ij}(t))$) of edge voltages which centre near 00 tend to produce higher dynamics, while those near the boundary of dynamic activity (i.e. boundary of the shaded region) produce less.
@@
However note that in Fig. A.1(b), although nonlinearity is maximal and most frequent within the specified region, there exists numerous highly nonlinear edges outside this region. The fact that nonlinearity can exist in edges which do not dynamically evolve implies the network connectivity has a significant effect on the overall nonlinear dynamics, and memristive edge dynamics is only a fraction of the cause and perpetuation of nonlinear effects. In fact, the network has the largest nonlinearity measure near $\alpha$ values with the most network induced nonlinearity (see Fig. A.4). Since only a small portion of memristive edges are evolving actively in Fig. A.1, the network appears to be underutilised; future studies (beyond simply changing the input layer) should optimise parameters which maximise the number of edges within this dynamically active region.
@@
It has been identified that both network and memristive dynamics causes nonlinearity. To investigate its exact origin, a simple 3-node toy model was constructed, with node 0 serving as the drain (ground) node, node 1 as readout and node 2 as the input electrode node (see schematic in Fig. A.2), which takes in a sine wave (shifted away from zero) as input. A highly nonlinear $x$ signal (extracted from the edge with the highest nonlinearity measure from Fig. A.1) is artificially injected into the $1-2$ edge. Both the modified and original $x_{1,2}$ values can be seen in Fig. A.2, where the modified $x_{1,2}$ display the classic behaviour of a memristive with high low-med dynamics measure as it never reaches the high conductance state. High frequency fluctuations can only be observed in the readouts with the modified $x$ instance, confirming these nonlinear features are directly caused by these types of memristive edge dynamics.
@@
Performing the same analysis on the 3-node network with a binary memristor model yields vastly different results. As seen in Fig. A.3, despite having the same modified $x_{1,2}$ signal as Fig. A.2, the high frequency fluctuations in the readouts are much more frequent, suggesting that the more physically realistic tunnelling memristor model may be more robust, and produce fewer high-frequency fluctuations when compared with the simpler binary model.
@@
Interestingly, the observed fluctuations are bounded between two values: the original sine wave readout and a constant zero signal (from the drain node). This is due to conductance paths (specifically $g_{0,1}$ and $g_{1,2}$) rapidly switching in amplitude across time, favouring one signal over the other at different times; in the actual network the high frequency fluctuations would be bounded by numerous other readouts in close proximity, as opposed to just two signals in this simple example. The gap at 34.5 Lyapunov times in Fig. A.3 reveals how these fluctuations disappear when $|x_{1,2}|>x_{\rm crit}$, which explains the absence of high frequency fluctuations with large low-high dynamics measures. Both Fig. A.2 and A.3 shows examples of how nonlinear responses in $x_{1,2}$ can cause similar behaviours in the neighbouring $x_{0,1}$ values, illustrating the spread of similar memristive dynamics within the network which has potential to cause observed nonlinear effects in the readouts.
@@ Appendix C Supporting Figures
Figs. A.4A.7 are referred to in the main text.
Figure A.4: The average nonlinearity measure $\bar{m}_{\rm nonlin}$ (defined in Eq. (20)) as a function of the scaling parameter $\alpha$. The simulations were performed on a neuromorphic network with $500$ nodes and $9,905$ edges. The dotted line at $\alpha=0.2$ volts indicate the default $\alpha$ used. Each data point was obtained from $100$ trials of the neuromorphic network simulation.
Figure A.5: Effect of node-node distances for input node selection on forecast times. Displayed are histograms of forecast times, with clustered inputs (each is within 2 distance of each other), and randomly selected inputs. Simulation performed on a neuromorphic network with 500 nodes and 9905 edges.
Figure A.6: Effect of memristor model on forecast times. Displayed are histograms of forecast times, with the (default) tunnelling, binary, and linear (i.e. resistor) models. Simulation performed on a neuromorphic network with $500$ nodes and 9905 edges, with $3,000$ data points per histogram.
Figure A.7: Effect of selected techniques (or lack thereof) used in this study on forecast times, displayed as histograms. “Modified bias” corresponds to the same parameters outlined in methods, with bias sampled from the interval $[-b,-b/2]\cup[b/2,b]$ “Uniform bias” has its values sampled from $[-b,b]$ (i.e. a uniform distribution $U(-b,b)$). “No bias” is equivalent to $\mathbf{b}_{\rm in}=0$. The skip connection is the additional $\mathbf{u}(t)$ term in Eq. (2). Simulation performed on a neuromorphic network with $500$ nodes and $9,905$ edges, with $3,000$ data points per histogram.

### v2

@@ IV Conclusions
These results validate the hypothesis that the internal and network dynamics of memristive networks can be used in a reservoir computing approach to learn complex chaotic dynamics. It was found that the network dynamics can be optimised for predicting the Lorenz63 attractor by appropriate scaling of the external input voltage to values that are not too low or too high. Optimal intermediate values of input voltage scaling were found to drive internal memristive states that traversed the entire dynamical range, from low to high conductance states, thereby maximising use of available internal degrees of freedom. Interestingly, this study also revealed that some nonlinear dynamical features, such as high–frequency fluctuations, were effectively filtered out by ridge regression in the external output layer. This suggests that physical reservoir computing may be robust against noise and variations due to the specific internal dynamics of the memristive material system. On the other hand, our results also indicate that it is possible to suppress such features by ensuring electrodes provide sufficient coverage of network nodes to sustain dynamic activity across the memristive edges.
@@ Appendix B Nonlinear Responses in Memristive Edges
Figure 17: (a) Total dynamics measure (i.e. $\delta_{\rm low,med}^{i,j}+\delta_{\rm med,high}^{i,j}+\delta_{\rm low,high}^{i,j}$) and (b) nonlinearity measure $m_{\rm nonlin}^{i,j}$ with respect to the maximum and minimum of each corresponding memristive edge voltage signal across time. The shaded regions are where edge voltage signal amplitudes can cross $V_{\rm set}$ and $V_{\rm reset}$, i.e. where $\dot{x}_{ij}$ may be nonzero. All data points are obtained from the same experiment for $\alpha=0.2\,{\rm V}$ in Fig. 11. Edges connected to the input nodes are omitted, and the figures are zoomed in for clarity.
@@
Figure 18: By artificially inducing a highly nonlinear (top, red) $x$ signal into the edge between node 2 and 1 of this simple 3-node toy model, unwanted nonlinear features (bottom, red) can be observed in the readouts of node 1, which are vastly different from the original (orange) signal. This also affect the $x$ (top, blue) signals of the edge between node 1 and 0. These memristive edges uses the tunnelling model.
@@
Figure 19: By artificially inducing a highly nonlinear (top, red) $x$ signal into the edge between node 2 and 1 of this simple 3-node toy model, unwanted nonlinear features (bottom, red) can be observed in the readouts of node 1, which are vastly different from the original (orange) signal. This also affect the $x$ (top, blue) signals of the edge between node 1 and 0. Unlike Fig. 18, these memristive edges uses the binary model.
@@
From Eq. (10) it is evident that $x$ will only increase in amplitude when the voltage amplitude at the corresponding memristive edge crosses $V_{\rm set}$, and decrease in amplitude when it crosses $V_{\rm reset}$. This will occur in two scenarios: if the maximum edge voltage signal (across time) of the edge between node $i$ and node $j$ satisfies $\max_{t}(v_{ij}(t))>V_{\rm set}$, while the minimum edge voltage satisfies $\min_{t}(v_{ij}(t))<V_{\rm reset}$, or $\max_{t}(v_{ij}(t))>-V_{\rm reset}$ and $\min_{t}(v_{ij}(t))<-V_{\rm set}$. This results in two overlapping regions where dynamical activity can be observed. Indeed as seen in Fig. 17(a), the dynamic measure is only nonzero within this specified (shaded) region, and it is clear that larger ranges (i.e. larger $\max_{t}(v_{ij}(t))$ and smaller $\min_{t}(v_{ij}(t))$) of edge voltages which centre near 0 tend to produce higher dynamics, while those near the boundary of dynamic activity (i.e. boundary of the shaded region) produce less.
@@
However note that in Fig. 17(b), although nonlinearity is maximal and most frequent within the specified region, there exists numerous highly nonlinear edges outside this region. The fact that nonlinearity can exist in edges which do not dynamically evolve implies the network connectivity has a significant effect on the overall nonlinear dynamics, and memristive edge dynamics is only a fraction of the cause and perpetuation of nonlinear effects. In fact, the network has the largest nonlinearity measure near $\alpha$ values with the most network induced nonlinearity (see Fig. 21). Since only a small portion of memristive edges are evolving actively in Fig. 17, the network appears to be underutilised; future studies (beyond simply changing the input layer) should optimise parameters which maximise the number of edges within this dynamically active region.
@@
It has been identified that both network and memristive dynamics causes nonlinearity. To investigate its exact origin, a simple 3-node toy model was constructed, with node 0 serving as the drain (ground) node, node 1 as readout and node 2 as the input electrode node (see schematic in Fig. 18), which takes in a sine wave (shifted away from zero) as input. A highly nonlinear $x$ signal (extracted from the edge with the highest nonlinearity measure from Fig. 17) is artificially injected into the 121-2 edge. Both the modified and original $x_{1,2}$ values can be seen in Fig. 18, where the modified $x_{1,2}$ display the classic behaviour of a memristive with high low-med dynamics measure as it never reaches the high conductance state. High–frequency fluctuations can only be observed in the readouts with the modified $x$ instance, confirming these nonlinear features are directly caused by these types of memristive edge dynamics.
@@
Performing the same analysis on the 3-node network with a binary memristor model yields vastly different results. As seen in Fig. 19, despite having the same modified $x_{1,2}$ signal as Fig. 18, the high–frequency fluctuations in the readouts are much more frequent, suggesting that the more physically realistic tunnelling memristor model may be more robust, and produce fewer high–frequency fluctuations when compared with the simpler binary model.
@@
Interestingly, the observed fluctuations are bounded between two values: the original sine wave readout and a constant zero signal (from the drain node). This is due to conductance paths (specifically $g_{0,1}$ and $g_{1,2}$) rapidly switching in amplitude across time, favouring one signal over the other at different times; in the actual network the high–frequency fluctuations would be bounded by numerous other readouts in close proximity, as opposed to just two signals in this simple example. The gap at 34.5 Lyapunov times in Fig. 19 reveals how these fluctuations disappear when $|x_{1,2}|>x_{\rm crit}$, which explains the absence of high–frequency fluctuations with large low–high dynamics measures. Both Fig. 18 and 19 shows examples of how nonlinear responses in $x_{1,2}$ can cause similar behaviours in the neighbouring $x_{0,1}$ values, illustrating the spread of similar memristive dynamics within the network which has potential to cause observed nonlinear effects in the readouts.
@@ Appendix C Supporting Figures
Figs. 2024 are referred to in the main text.
Figure 20: Degree distribution of the two physical neuromorphic networks simulated in this study. Each marker represents the fraction (relative frequency) of each degree. The curves show the best fit skewed normal distribution, i.e. the networks have an approximately skewed log-normal degree distribution. For the $2,000$-node network, the shape, location and scale parameters of the best fit distribution are 9.0-9.0, $4.3$ and $1.0$ respectively. For the 500-node network, the parameters are 3.7-3.7, $4.5$ and $1.1$ respectively.
Figure 21: The average nonlinearity measure $\bar{m}_{\rm nonlin}$ (defined in Eq. (20)) as a function of the scaling parameter $\alpha$. The simulations were performed on a neuromorphic network with $500$ nodes and $9,905$ edges. The dotted line at $\alpha=0.2$ volts indicate the default $\alpha$ used. Each data point was obtained from $100$ trials of the neuromorphic network simulation.
Figure 22: Effect of node-node distances for input node selection on forecast times. Displayed are histograms of forecast times, with clustered inputs (each is within 2 distance of each other), and randomly selected inputs. Simulation performed on a neuromorphic network with 500 nodes and 9905 edges.
Figure 23: Effect of memristor model on forecast times. Displayed are histograms of forecast times, with the (default) tunnelling, binary, and linear (i.e. resistor) models. Simulation performed on a neuromorphic network with $500$ nodes and 9905 edges, with $3,000$ data points per histogram.
Figure 24: Effect of selected techniques (or lack thereof) used in this study on forecast times, displayed as histograms. “Modified bias” corresponds to the same parameters outlined in methods, with bias sampled from the interval $b,-b/2]\cup[b/2,b]$ . “Uniform bias” has its values sampled from [b,b][-b,b] (i.e. a uniform distribution U(b,b)U(-b,b)). “No bias” is equivalent to $\mathbf{b}_{\rm in}=0$. The skip connection is the additional $\mathbf{u}(t)$ term in Eq. (2). Simulation performed on a neuromorphic network with $500$ nodes and $9,905$ edges, with $3,000$ data points per histogram.